OpenStudy (anonymous):

if you count by 25s from 575,what will be the 12th number?

OpenStudy (anonymous):

nth term * rate +base 12 25 575

OpenStudy (anonymous):

Tn=a+(n-1)d where a is the first term n is the required one d is the difference between the terms so...Tn=575+(11)25 Tn=850